Let a+b+c+ab+bc+ca+abc  = 1000


Then The Possible Value Of a+b+c Is 


A) 28


B) 36


C) 39 


D) 42

Grade:9

1 Answers

jitender lakhanpal
11 years ago

hi prakhar

the expression is (a+1)(b+1)(c+1) - 1 = 1000

(a+1)(b+1)(c+1)=1001 = 7*11*13

from here we can conclude a = 6, b= 10 , c= 12

can be the possible values

We are  here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

 

jitender ''

askiitians expert